# ATOMIUM: Dust and tracers of binarity in the continua

### v1

@@ 3 Continuum imaging
The combined visibility data provide sensitivity to emission on scales from a few tens of mas up to 8–10″. However, the visibility plane coverage is not complete nor evenly distributed. To avoid artefacts due to missing spacings between the longest baselines, we applied a slight taper in imaging to weight down their contribution, providing resolutions 50 mas. Residual calibration errors and antenna position errors limit the dynamic range555Dynamic range is (peak flux) / (rms). to about 1000, so for the brightest stars, e.g. R Hya, symmetric negative and positive artefacts may be present at % of the peak. Tables 5, 6, 7, and 8 include the dynamic ranges for for the extended, mid, compact and combined data, respectively. The ACA continua have relatively low dynamic ranges, from 4 for RW Sco to 50 for T Mic (Table 9).
@@
Smooth emission on scales much greater than 10″ will not be detected, but any extended dust on around this scale may cause bowl-like artefacts due to being detected by just a few baselines. Artefacts may also arise from intermediate under-sampled scales. In general, for the brightest stars, if there are regions of negative flux (indicated by dotted contours at the 3 and levels), corresponding positive regions at those levels or weaker should be treated with caution. The noise rms given in Table 2, used to estimate the accuracy of quantities derived from UD fits, are measured close to the stellar position. The background noise far from the star is given in Tables 5–8 for the extended, mid, compact and combined array data respectively. The combined noise at ¿ 1″ is mJy.
@@
The continuum images for the extended, mid and combined data are plotted, per star, in Figures 2–LABEL:fig:ahsco, as indicated for each star in Table 1. We also include in these figures combined images of the residuals after subtracting uniform discs representing the AGB stars and sometimes including dust close to the stars (as described in Sect. 2.4). The compact array data are plotted for all observed stars in Fig. 11 and the Band 3 ACA data are shown in Fig. 12. In general, the compact configuration images are unresolved or marginally resolved, with some stars exhibiting small ( beam) irregular extended regions of flux detected at the level. We can determine that the bulk of the flux is unresolved because it mimics the elliptical shape of the beam down to the 5 or level. All of the ACA images are unresolved, owing to the large beams (see Table 9). The only unusual ACA result is for RW Sco and this is discussed further in Appendix C.
In Fig. 13, we plot the visibility amplitudes of the continuum data for each star against the uv distance and baseline length (the bottom and top axes). These plots also show the UD fits (black curves, see Sect. 2.4). For W Aql, U Her, R Hya, S Pav, IRC+10011, and AH Sco, these plots show excesses at short baselines, indicating the presence of large scale structures. This is also reflected in the imaging, as discussed in Sect. 3.2 for U Her, R Hya, S Pav, and IRC+10011 and in Sections 3.4.2 and 3.4.5 for W Aql and AH Sco, respectively.
@@
In the following sections we discuss some stars grouped by similar continuum features and some stars individually when they exhibit relatively unique features. These groupings are summarised in Table LABEL:tab:class.
3.1 Spatially compact continuum images
Several stars in our sample (R Aql, U Del, V PsA, RW Sco, KW Sgr) have relatively featureless continua, with no significant () deviations from circular symmetry or the shapes of the beams for the extended, combined or UD-subtracted images. In the case of U Del, the compact image (Fig. 11) exhibits only very minor deviations from elliptical at the level. The UD subtracted image does not exhibit a peak at the position of the star and the only features present in the whole image are at levels of – but smaller than the beam, making them non-significant (Fig. LABEL:fig:udel). The compact continuum of V PsA has narrow plumes at a level of extending to the northeast and southwest (Fig. 11), but these features are narrower than the beam and hence uncertain. The UD subtracted image of V PsA (Fig. LABEL:fig:vpsa) retains a peak at the position of the star, indicating significant emission in this region, not counted for by the UD fit, unlike the case of U Del. The continuum maps of RW Sco are similarly featureless, with some emission present at the continuum peak position after the UD has been subtracted (Fig. LABEL:fig:rwsco). However, we note that the UD fit to RW Sco was very uncertain, meaning that the UD subtracted image is also very uncertain.
The extended and combined continuum maps of R Aql, plotted in Fig. 2, have very high dynamic ranges, in excess of 2000, resulting in increased uncertainty such that features in the map of the peak should be treated with caution, especially if symmetric. The extended and combined continuum images of R Aql are relatively featureless, but the mid image shows some irregular extended emission around the peak (Fig. 2), possibly indicating dusty regions offset from the stellar position. Some similar extended emission is also seen in the compact continuum image (Fig. 11). The UD subtracted image shows a significant amount of flux centred on the continuum peak, in addition to fainter regions of flux at angular separations of 0.2–0.4″ from the star.
KW Sgr, a RSG not observed with the compact configuration of ALMA, has a highly elliptical beam for the mid data and some small asymmetries in the continuum at levels of but generally smaller than the area of the beam, and hence not significant. The UD subtracted image (Fig. LABEL:fig:kwsgr) shows some extended dust to the east south-east of the continuum peak, with a confidence of , and a little more emission at the position of the continuum peak. This is discussed further in Sect. 5.1.6.
3.2 Stars with significant extended flux
@@
Significant extended flux is seen in the combined and UD-subtracted continuum images of U Her (Fig. LABEL:fig:uher), R Hya (Fig. LABEL:fig:rhya), S Pav (Fig. LABEL:fig:spav), and IRC+10011 (Fig. LABEL:fig:irc+10011). In these cases, the flux is relatively symmetric about the stellar positions, though not perfectly so. U Her and R Hya also exhibit some extended emission in their mid continuum maps.
@@ 3.2.1 U Her
The continuum maps of U Her in Fig. LABEL:fig:uher show significant extended emission surrounding the continuum peak, most clearly apparent in the UD-subtracted image. Even if we disregard the contour, there is still significant excess flux. The exception is the compact map (Fig. 11), which is featureless aside from a detached region to the west-northwest at around and slightly smaller than the size of the synthetic beam, hence not significant. There are several protrusions detected at levels of surrounding the continuum peak in the extended map. These are asymmetric, so unlikely to arise from the data reduction. However, there are two negative regions (down to ) which may indicate regions of resolved-out flux.
The combined continuum map shows larger regions of extended emission, more clearly visible in the UD-subtracted image. The larger-scale emission at extends to separations of around 0.4″ to the south, west and north, but only out to to the east-northeast. The emission is not distributed evenly and is clumpier to the north and northwest. In the extended, mid and combined maps, there is a small region (at different scales for each map) with less emission to the south of the continuum peak. The cause of this region of lower emission is unclear, but the fact that we see it at all scales suggests that either a continuous or repeating process is causing it.
@@ 3.2.2 R Hya
The continuum observations of R Hya (Fig. LABEL:fig:rhya) have the highest dynamic range of any star in our sample, exceeding 1000 for all individual array configurations, and reaching 6600 for the combined data (see Tables 5–8). This may be partly owing to the fact that R Hya is the closest star in our sample (distance = 126 pc). The roughly symmetric negative regions surrounding the continuum peak in the extended and combined maps shown in Fig. LABEL:fig:rhya are the result of amplitude errors. However, the UD-subtracted map shows significant extended flux at levels of out to 0.15–0.3″ from the stellar position.
The mid continuum map has scattered regions of emission around the central continuum peak. These include emission detected with a certainty of extended out to . The compact map is only marginally resolved and relatively featureless (Fig. 11). It is also the only compact map with a dynamic range .
@@ 3.2.4 IRC+10011 (WX Psc)
In the compact continuum map (Fig. 11) we find two small regions of flux detected to the northwest and southeast of the continuum peak. These regions are only detected at a level of , and with the contours smaller than the synthetic beam, making them unreliable detections. Checking a UD-subtracted combined image made with a coarse taper and comparable synthetic beam to the compact data does not recover these features. Hence we must conclude they are not real. This coarse taper combined image does recover an extension to the north, increasing our confidence that this feature is real.
3.3 Bipolar continuum features
Identifying bipolar structures in the continuum maps can be challenging, as any symmetric features (of positive or negative flux) may be the result of amplitude errors. These are more likely to be present for higher dynamic ranges. Since our sample contains many bright sources, we must be careful when interpreting any symmetric bipolar features. With this in mind, we have identified a group of stars which exhibit some bipolar features. These are IRC10529, T Mic, and VX Sgr.
@@ 3.3.1 IRC10529 (V1300 Aql)
The extended map of IRC10529 (Fig. LABEL:fig:irc-10529) features some small symmetric protrusions to the northwest and southeast, which may be amplitude errors rather than real features. However, there is a small region of extended flux to the southwest which is likely real and is not mirrored on the other side. The mid image, suffering from less resolved out flux, also reveals a region of extended flux to the southwest. This extended flux is clearer in the combined map where we can see some emission to the west-southwest and east-northeast. The UD-subtracted map, which emphasises lower surface brightness features thanks to its lower angular resolution, highlights a band of flux running along an axis east from north. There are also some perpendicular extensions which are not symmetric about the location of the continuum peak. A line along the same axis ( east from north) passes between the four outer “points” of the UD-subtracted map (e.g. assuming a similar shape to the Red Rectangle protoplanetary nebula, Cohen et al. 2004, but on a much smaller scale).
@@ 3.3.2 T Mic
The mid and combined maps of T Mic (Fig. LABEL:fig:tmic) are relatively featureless. There are some regions of symmetric low-level flux in the extended map, but this may be the result of amplitude errors rather than real features. A few irregular regions, smaller than the beam size, are present on the compact map (Fig. 11), but their significance is not certain. The most interesting result for T Mic is the UD subtracted image. Here, a region to the northwest is seen at a significance of 3– and a more elongated region at around is seen opposite, to the southeast. The coarser beam size of the UD-subtracted image increases the signal to noise of features that are not confidently detected in the combined map with a smaller restoring beam. Drawing a line between the continuum peak and the northern feature in the UD map, we find that it lies approximately west of north. This is in the same direction as the brightest feature seen in polarised light in the SPHERE map of Montargès et al. (2023), raising confidence in this feature. Although this line also passes through some of the southern features, we find that the southern features lie within a cone covering 20– east of south. The best agreement for both northern and southern features is described by a line running west of north, which is also in reasonable agreement with the (uncertain) northern and southern features on the extended map.
@@ 3.3.3 VX Sgr
Some asymmetric features are seen for all the continuum maps of VX Sgr (Fig. LABEL:fig:vxsgr, 11). These are most pronounced for the extended and combined maps, where emission is seen predominantly to the south and north of the continuum peak. Once again, this emission is most pronounced in the UD subtracted map. Considering only the emission closest to the location of the continuum peak (e.g. the emission enclosed by the contour), we find that the bipolar axis is close to running due north-south, with an uncertainty of only a few degrees ().
3.4 Unique continuum features
@@ 3.4.1 GY Aql
GY Aql is unusual in that the continuum maps from the extended and compact array configurations are relatively featureless (Fig. 2, Fig. 11), but a strongly detected feature is seen to the southeast in the mid and combined maps. In the mid map, this feature appears attached to the continuum peak and is detected at a level . When constructing the combined image, we found that choosing a restoring beam that was too fine would result in this feature having a very low surface brightness and not being detected above the noise, beyond a small region with the brightest flux (visible at an offset of 0.5, in Fig. 2). Hence, we chose a coarser resolution for the combined map than was used for most other stars (see Table 8) to emphasise this feature, which appears detached and which we will refer to as a “bar”, though it may be part of a spiral, since spiral-like structures are seen in the CO emission (Decin et al., 2020). As can be seen in the UD subtracted map, the continuum flux mainly originates from this bar and from regions close to the star, likely indicating a significant concentration of dust in this bar.

### v2

@@ 3 Continuum imaging
The combined visibility data provide sensitivity to emission on scales from a few tens of mas up to 8–10″. However, the visibility plane coverage is not complete nor evenly distributed. To avoid artefacts due to missing spacings between the longest baselines, we applied a slight taper in imaging to weight down their contribution, providing resolutions 50 mas. Residual calibration errors and antenna position errors limit the dynamic range555Dynamic range is (peak flux) / (rms). to about 1000, so for the brightest stars, e.g. R Hya (Fig. LABEL:fig:rhya), symmetric negative and positive artefacts may be present at % of the peak. Tables 7, 8, 9, and 10 include the dynamic ranges for for the extended, mid, compact and combined data, respectively. The ACA continua have relatively low dynamic ranges, from 4 for RW Sco to 50 for T Mic (Table 11 and Fig. 15).
@@
Smooth emission on scales much greater than 10″ will not be detected, but any extended dust on around this scale may cause bowl-like artefacts due to being detected by just a few baselines. Artefacts may also arise from intermediate under-sampled scales. In general, for the brightest stars, if there are regions of negative flux (indicated by dotted contours at the 3 and levels), corresponding positive regions at those levels or weaker should be treated with caution. The noise rms given in Table 2, used to estimate the accuracy of quantities derived from UD fits, are measured close to the stellar position. The background noise far from the star is given in Tables 7–10 for the extended, mid, compact and combined array data respectively. The combined noise at ¿ 1″ is mJy.
@@
The continuum images for the extended, mid and combined data are plotted, per star, in Figures 2–LABEL:fig:ahsco, as indicated for each star in Table 1. We also include in these figures combined images of the residuals after subtracting uniform discs representing the AGB stars and sometimes including dust close to the stars (as described in Sect. 2.4). The compact array data are plotted for all observed stars in Fig. 14 and the Band 3 ACA data are shown in Fig. 15. In general, the compact configuration images are unresolved or marginally resolved, with some stars exhibiting small ( beam) irregular extended regions of flux detected at the level. We can determine that the bulk of the flux is unresolved because it mimics the elliptical shape of the beam down to the 5 or level. All of the ACA images are unresolved, owing to the large beams (see Table 11).
@@
In the following subsection we discuss the issue of resolved out flux. Subsequently we discuss the unresolved continuum emission, then some stars grouped by similar continuum features and some stars individually when they exhibit relatively unique features. These groupings are summarised in Table LABEL:tab:class.
3.1 Resolved-out flux
As noted in Sect. 2.2, ALMA is not guaranteed to recover all the flux associated with a source and may resolve out some smooth large scale flux. We are able to check the impact of this effect for four ATOMIUM stars which are also in the DEATHSTAR survey (Ramstedt et al., 2020). DEATHSTAR observed a large sample of AGB stars using the ALMA ACA in Bands 6 and 7. For their Band 6 observations, covering similar frequencies to ATOMIUM albeit over a smaller range, the MRS was (Ramstedt et al., 2020). We compared the self-calibrated Band 6 continuum fluxes from DEATHSTAR with those from our ATOMIUM combined and compact continuum images, for the overlapping stars (IRC10529, SV Aqr, T Mic and IRC+10011). We found no evidence of more flux being resolved out for the ATOMIUM data than for the DEATHSTAR data.
However, this does not mean that all the flux has necessarily been recovered. Herschel/PACS images of dust at 70 and 160 m are available for some of the stars in our sample. For W Aql, Gru, R Hya, and T Mic, extended flux is seen on scales of from the position of the star, indicating significant dust at larger scales than can be recovered with ALMA (Cox et al., 2012). This is significant because at longer wavelengths (i.e. 1.2–1.3 mm for ALMA Band 6) we might detect cooler dust, which we would expect to be located further from the star. However, since ALMA cannot recover flux on such large scales — scales that can be larger than the ALMA field of view for a single pointing — we would not expect the majority of flux from such a large dust shell to be recovered, even when using the ACA. Comparing our total combined fluxes (Table 10) with single-dish observations from Altenhoff et al. (1994) and Dehaes et al. (2007), we can estimate the recovered fluxes for R Aql, R Hya, and IRC+10011 as around 60%, 70%, and 6–12%, respectively. The large amount of resolved out flux for IRC+10011 (88–94% lost) fits with the negative artefacts we see in Fig. LABEL:fig:irc+10011.
To give an indication of what fraction of flux comes from regions offset from the continuum peak, we compared the UD flux density (see Table 2 and Sect. 2.4) with the total flux recovered in our combined images (Table 10). Overall, we found no significant correlation with distance to the stars. Excluding the RSGs, we found a weak positive correlation between the flux ratio and distance, which is expected, since a more distant source will appear smaller on the sky and hence be less susceptible to resolved-out flux. Also, for the closest sources, the full extent of the circumstellar envelope may be larger than the primary beam of the ALMA 12-m array. Across our sample, there is a large spread in the ratios of total flux to UD flux, from 1.07 (U Del) to 2.24 (IRC10529). U Del having the lowest flux ratio fits with what we found from the UD-subtracted map (Fig. LABEL:fig:udel) which shows that the majority of the continuum flux is associated with a uniform disc of emission at the stellar position and that there is essentially no extended flux recovered by ALMA. The stars with the highest ratio (IRC10529 followed by W Aql) exhibit flux away from the continuum peak. Some of the other stars with lower flux ratios nevertheless have extended features, likely arising from dust, in their continuum maps, or are known to be dusty based on other observations (e.g. Herschel/PACS imaging Cox et al., 2012; Maercker et al., 2022). This discrepancy could be caused by resolved-out flux, amplitude errors (especially for sources with high dynamic ranges ) or a combination of factors.
We also checked whether there was any correlation between the ratio of total flux and UD flux density with the apparent or physical UD radii and found no trends. We also found no correlation between the total flux or UD flux with mass-loss rate, distance, or mass-loss rate scaled by distance.
3.2 Resolution-limited continuum images
In general, we expect spatially resolved emission to appear larger than the restoring beam of the observing array. For emission with spatial extents much larger than the beam, this is trivially apparent (see for example the ATOMIUM CO emission, Decin et al. 2020). However, a bright point-like source with a high dynamic range can appear to be larger in the image plane than the reported beam size, because of the Gaussian nature of the beam. This is explained in detail in Appendix F, where we lay out criteria for checking whether bright sources show evidence of resolved emission. Note that fitting in the visibility plane, as we did for the UD fits (Sect. 2.4), avoids apparent extension due to convolution with the beam, but is still S/N limited.
The most sensitive continuum images, capable of revealing the most extreme small- and large-scale features, were made from the combined data. Most of our sources exhibit some extended emission in the combined maps, though in some cases, such as KW Sgr (Fig. LABEL:fig:kwsgr), the extended emission is not apparent until we subtract a UD from the data and lower the resolution to increase sensitivity (Sect. 2.4). For the particular example of KW Sgr, this is discussed further in Sect. 5.1.6. Two stars, U Del (Fig. LABEL:fig:udel) and RW Sco (Fig. LABEL:fig:rwsco), do not show any extended emission in the combined maps or the UD-subtracted maps. U Del is so well-described by the UD that there is no significant flux in the UD subtracted map at all. T Mic (Fig. LABEL:fig:tmic) and V PsA (Fig. LABEL:fig:vpsa) appear unresolved in the combined maps, but the UD subtracted maps reveal some small (comparable to the beam) regions of emission offset from the continuum peaks. The combined continuum maps for the remaining stars all show more significant extended emission and/or asymmetries, which are described in the following subsections.
When considering the individual array images in the context of our criteria for resolution, we find that almost all of our compact images (Fig. 14) are close to the resolution criterion defined in Appendix F, i.e. within 10% at the level, with only a few stars such as W Aql and S Pav showing evidence of extended emission of the resolution criterion. For observations with the mid configuration, a larger portion of the images show evidence of (partially) resolved extended emission. Some cases of faint extended emission seem to be below the sensitivity limit of the mid data, but above in the combined data, making sources such as S Pav (Fig. LABEL:fig:spav) appear unresolved in the mid images. Most images taken with the extended array show evidence of resolved out flux to varying degrees, with only a few stars for which clumpy extended emission is recovered. A prime example of this is Gru (Fig. LABEL:fig:pi1gru), which we discuss in detail in Sect. 3.5.5.
3.3 Stars with relatively uniform extended flux
@@
Based on the criteria laid out in Appendix F, the majority of the stars in our sample have some extended continuum emission associated with them. Of these, relatively uniform extended flux is seen in the combined and UD-subtracted continuum images of U Her (Fig. LABEL:fig:uher), R Hya (Fig. LABEL:fig:rhya), S Pav (Fig. LABEL:fig:spav), and IRC+10011 (Fig. LABEL:fig:irc+10011). In these cases, the flux is relatively symmetric about the stellar positions, though not perfectly so. U Her and R Hya also exhibit some extended emission in their mid continuum maps.
@@ 3.3.1 U Her
The continuum maps of U Her in Fig. LABEL:fig:uher show significant extended emission surrounding the continuum peak, most clearly apparent in the UD-subtracted image, where the regions of extended emission much larger than the beam. The mid and extended maps show asymmetric features at the level.
In the combined map, the larger-scale emission at extends to separations of around 0.4″ to the south, west and north, but only out to to the east-northeast. The emission is not distributed evenly and is clumpier to the north and northwest. In the extended, mid and combined maps, there is a small region (at different scales for each map) with less emission to the south of the continuum peak. The cause of this region of lower emission is unclear, but the fact that we see it at all scales suggests that either a continuous or repeating process is causing it.
@@ 3.3.2 R Hya
The continuum observations of R Hya (Fig. LABEL:fig:rhya) have the highest dynamic range of any star in our sample, exceeding 1000 for all individual array configurations, and reaching 6600 for the combined data (see Tables 7–10). This may be partly owing to the fact that R Hya is the closest star in our sample (distance = 126 pc). The roughly symmetric negative regions surrounding the continuum peak in the extended and combined maps shown in Fig. LABEL:fig:rhya are the result of amplitude errors. However, the UD-subtracted map shows significant extended flux at levels of out to 0.15–0.3″ from the stellar position.
The mid continuum map has scattered regions of emission around the central continuum peak. These include emission detected with a certainty of extended out to .
@@ 3.3.4 IRC+10011 (WX Psc)
3.4 Elongated continuum features
Identifying elongated structures in the continuum maps can be challenging, as any symmetric features (of positive or negative flux) may be the result of amplitude errors. These are more likely to be present for higher dynamic ranges. Since our sample contains many bright sources, we must be careful when interpreting any symmetric elongated features. With this in mind, we have tentatively identified a group of stars which exhibit some elongated features. These are IRC10529 (Fig. LABEL:fig:irc-10529), T Mic (Fig. LABEL:fig:tmic), and VX Sgr (Fig. LABEL:fig:vxsgr).
@@ 3.4.1 IRC10529 (V1300 Aql)
The extended map of IRC10529 (Fig. LABEL:fig:irc-10529) features some small symmetric protrusions to the northwest and southeast, which may be amplitude errors rather than real features. However, there is a small region of extended flux to the southwest, in the extended and mid maps, which is likely real and is not mirrored on the other side. The combined map shows an elongated emission structure, with some emission extending from the continuum peak to the west-southwest and east-northeast. This elongated structure, which runs along an axis east from north, is detected with more confidence in the UD-subtracted map, which emphasises lower surface brightness features thanks to its lower angular resolution. This axis is comparable to the axis about which the positive and negative velocities are separated in the SiO moment 1 map shown in Decin et al.’s (2020) Figure S60. There are also some perpendicular extensions which are not symmetric about the location of the continuum peak, and which are somewhat reminiscent of the shape to the Red Rectangle protoplanetary nebula (Cohen et al., 2004), but on a much smaller scale.
@@ 3.4.2 T Mic
Although the individual array images of T Mic (Fig. LABEL:fig:tmic) are unresolved, the UD subtracted image reveals a region to the northwest at a significance of 3– and a more elongated region, at around , to the southeast. The coarser beam size of the UD-subtracted image increases the signal to noise of features that are not confidently detected in the combined map with a smaller restoring beam. Drawing a line between the continuum peak and the northern feature in the UD map, we find that it lies approximately west of north. This is in the same direction as the brightest feature seen in polarised light in the SPHERE map of Montargès et al. (2023), as we plot in Fig. 3, raising confidence that this elongated structure is real. The angle of this feature is approximately perpendicular to the angle separating positive and negative velocities in the 30SiO () moment 1 map presented in Decin et al. (2020).
@@ 3.4.3 VX Sgr
Some asymmetric features are seen for all the continuum maps of VX Sgr (Fig. LABEL:fig:vxsgr). These are most pronounced for the extended and combined maps, where emission is seen predominantly to the south and north of the continuum peak. Once again, this emission is most pronounced in the UD subtracted map. Considering only the emission closest to the location of the continuum peak (e.g. the emission enclosed by the contour), we find that the elongation axis is close to running due north-south, with an uncertainty of only a few degrees (). We note that this emission could be the result of episodic mass loss, as discussed in Sect. 5.1.7.
3.5 Highly asymmetric continuum features
@@ 3.5.1 GY Aql
GY Aql is unusual in that the continuum maps from the extended and compact array configurations show unresolved emission (Fig. 2, Fig. 14), but a strongly detected feature is seen to the southeast in the mid and combined maps. In the mid map, this feature appears attached to the continuum peak and is detected at a level . When constructing the combined image, we found that choosing a restoring beam that was too fine would result in this feature having a very low surface brightness and not being detected above the noise, beyond a small region with the brightest flux (visible at an offset of 0.5, in Fig. 2). Hence, we chose a coarser resolution for the combined map than was used for most other stars (see Table 10) to emphasise this feature, which appears detached and which we will refer to as a “bar”, though it may be part of a spiral, since spiral-like structures are seen in the CO emission (Decin et al., 2020). As can be seen in the UD subtracted map, the continuum flux mainly originates from this bar and from regions close to the star, likely indicating a significant concentration of dust in this bar.
3.5.2 R Aql
In constructing the combined continuum image, we found suspiciously regular positive and negative patches of emission, larger than the synthesised beam, at levels of around the continuum peak. This suggests resolved-out emission. Tapering to give more weight to shorter baselines, i.e. coarsening the resolution, alleviated some of this effect and allowed us to recover more flux and structures in the combined map, which is shown in Fig. 2. The extended continuum emission is patchy and irregularly distributed around the continuum peak. Hints of these structures can also be seen in the marginally resolved mid continuum map. The extended continuum map of R Aql has a very high dynamic ranges, in excess of 2000, resulting in increased uncertainty such that features in the map of the peak should be treated with caution, especially if symmetric. All the larger scale structures visible in the mid and combined maps are resolved-out in the extended map.